OpenStudy (anonymous):

x³-3x+4 How do you Descartes' Rule and the Fundamental Theorem of Algebra to determine the zeros(complex and real)?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

descartes i believe looks at sign changes for f(x) and f(-x) to determine the number of possible roots; but it doesnt say anythiing about the value of those roots

OpenStudy (amistre64):

the fundamental thrm of algebra simply says that a poly has as many roots as its degree, but it doesnt say what the value of those roots are either

OpenStudy (anonymous):

I know Descartes' Rule.. and f(x) sign changes tells me the maximum numbers of positive roots. f(-x) tells me the max number of negative roots.

OpenStudy (anonymous):

But I don't understand Fund. The. of Alg. But, I have to write about how to find complex roots using those two terms.

OpenStudy (amistre64):

and those maxs can be altered by 2s to determine complex roots

OpenStudy (anonymous):

I don't know how to find complex roots.

OpenStudy (amistre64):

spose you find 3 sign changes in f(x); then there are at most 3 positive roots, or at least 1 positive root if we consider complexes; 3real - 2complex = 1 real

OpenStudy (anonymous):

Okay, so this equation will have only 2 complex roots?

OpenStudy (amistre64):

its not a gaurentee, just a means of considering possibilities

OpenStudy (anonymous):

Okay. How do I know what the roots are for this polynomial?

OpenStudy (amistre64):

x^3 - 3x + 4 ^ ^ 2 sign changes; max of 2 positive roots, or 0 positive roots 2 real roots - 2 complex roots = 0 real roots

OpenStudy (amistre64):

the only way to find the roots is to work thru the problem using other methods not descrbed in the context of the question

OpenStudy (anonymous):

Oh.. well what method would I use to solve for complex roots?

OpenStudy (amistre64):

we can use the rational root thrm to determine of it has ANY real roots, and that would allow us to factor the poly into more workable forms

OpenStudy (anonymous):

Mk, how do you do that?

OpenStudy (anonymous):

I'm sorry... I'm lost.

OpenStudy (amistre64):

its a modified trial and error process :) you narrow down the option to a set of choices that have the best chance of being a real root; if it is none of those options, then it has no real roots

OpenStudy (anonymous):

So, should I try -1,1,-4, and 4?

OpenStudy (amistre64):

\[\pm\frac{factors~last\#}{factors~first\#}\]

OpenStudy (amistre64):

hmm: i would try +- 1,2,4

OpenStudy (anonymous):

Oh, I forgot two. :P and okay, after I figure out which works, how do I determine the complex roots?

OpenStudy (amistre64):

with any luck, it reduces to a prime poly that can be manipulated im not sure, but we might be able to include complex factors of the first and last to add to the options list; but ive never really tried to do that

OpenStudy (amistre64):

lets try out our reals :)

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

Can use synthetic division?

OpenStudy (amistre64):

x^3-3x+4 ; x=1 1-3+4 not = 0 x=2 8-6+4; not 0 x=4 64-12+4; not 0 no positive real roots; try the negatives

OpenStudy (amistre64):

we could use synthetic if need be, but a calculator works fine if you can use one :)

OpenStudy (anonymous):

None of them work.

OpenStudy (anonymous):

The negatives.

OpenStudy (amistre64):

so it doesnt have rational roots; but it has to have at least on real root (its an odd degree) .... so it would have to be an irrational number

OpenStudy (anonymous):

Okay, do I just figure that out by guess & check? Or is there a way I can figure out the number by looking at the equation?

OpenStudy (amistre64):

there is no simple way to do this by hand http://www.wolframalpha.com/input/?i=x%5E3-3x%2B4 notice the complicated nature of the root

OpenStudy (amistre64):

there are calculus methods that can get us closer and closer if we required a approximate value to a certain decimal accuracy

OpenStudy (anonymous):

I see... I am allowed to make up my own polynomial.. So could you make an easier one for me to figure out the roots for?

OpenStudy (amistre64):

its easier to manufacture it from roots, that way you know you have something workably simple :) how many roots would you like, how many complex and how many neg/pos ?

OpenStudy (anonymous):

Nothing that causes tooo much work to explain in my essay. Something simple that still has one complex root and it's conjugate + a few real roots. Hehe :)

OpenStudy (amistre64):

lets use roots: -3,-1,2, 3i, -3i \[(x+3)(x+1)(x-2)(x-3i)(x+3i)\]\[\hspace{3em} = x^5 +2x^4+4x^3+12x^2-45x-54\]

OpenStudy (anonymous):

That was a little backwards from what I am supposed to do. Hehe. But it works! Now I have to explain how I used Descartes' Rule and Fund. Theo. of. Alg to find those zeros. I'll just restate the definition, I guess? & add other things like Synthetic Division and whatnot.

OpenStudy (amistre64):

that would be fine :)

OpenStudy (anonymous):

Thanks for the help :)

OpenStudy (amistre64):

yw, good luck ;)

OpenStudy (anonymous):

Can you give me one more example like the one you just gave me?

OpenStudy (amistre64):

its the same process, define a few roots, and multiply them together to get the poly

OpenStudy (anonymous):

Ohhhhh, okay. So it's really anything I want it to be?

OpenStudy (amistre64):

yep, anything you can dream of roots: -1,1,+- i (x+1)(x-1)(x+i)(x-i) = x^4-1